OpenStudy (anonymous):

Rock music accounted for 34.5% of the total sales of recorded music in 1995, and 22.8% of the total sales of recorded music in 2005. Find the average rate of change of the percentage of sales of rock music from 1995 to 2005. Round to the nearest hundredth of a percent.

OpenStudy (anonymous):

Hi there. Any ideas?

OpenStudy (anonymous):

Do I just subtract 22.8% from 34.5% and divide the answer by 10?

OpenStudy (anonymous):

Yes, slope = Δy / Δx

OpenStudy (anonymous):

Wait....I just did that....and got 1.17, which is apparently wrong.

OpenStudy (anonymous):

-11.7% / 10 = -.0117

OpenStudy (anonymous):

Pay attention of decreasing percentage!

OpenStudy (anonymous):

Wait...no!

OpenStudy (anonymous):

That answer's apparently wrong too!

OpenStudy (anonymous):

(1995, 34.5% ) and ( 2005, 22.8%) slope = Δy / Δx = ( 34.5% - 22.8%) / ( 1995 - 2005) = - 11.7% / 10 = - 1.17%

OpenStudy (anonymous):

Do you know keep the percentage or not?

OpenStudy (anonymous):

Okay..got it. No, I'm afraid not.

OpenStudy (anonymous):

Just give the answer -1.17% = -.0117

OpenStudy (anonymous):

Alright. Thank you a tremendous amount, Chlorophyll!

OpenStudy (anonymous):

Since you had showed you know the calculation, I didn't go into details at first :)